OpenStudy (anonymous):

What is the cube root of 512m^12n^15

OpenStudy (welshfella):

cube root of m^12 = (m^12)^1/3 = m^4

OpenStudy (welshfella):

cube root of n^15 = (n^15)^1/3 = n^(15*1/3) = n^5

OpenStudy (anonymous):

thanks

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

cube root of 512

OpenStudy (welshfella):

8 *8*8 = 64*8= 512

OpenStudy (anonymous):

preciate that